# Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

### v1

@@ 5.2 Fine-tuning on Downstream Tasks
We further evaluated SelectiveMAE across three key downstream tasks: scene classification, object detection, and semantic segmentation. In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[82], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models. These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning based approaches [83]. Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE). To ensure fairness, all experiments followed to consistent fine-tuning settings. Supplementary materials and code provide comprehensive experimental details.
@@
Scene Classification. We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders. We leverage two scene classification datasets: AID [84] and RESISC-45 [77]. Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code. Evaluation is based on overall accuracy (OA). Table 4 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline.
Besides, it outperforms RVSA and others when scaled to ViT-L. This demonstrates that, with our SelectiveMAE training method on the RS-4M dataset, the model can learn strong feature representations and scale efficiently.
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [85] and its improved variant DIOR-R for oriented object detection [86], both comprising RGB images. Following the methodologies of previous works [18, 37], we maintained consistent experimental setups, employing Faster-RCNN [87] and Oriented-RCNN [88] as detectors for each dataset. Detailed experimental configurations are available in the accompanying code. The results are summarized in Table 4. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37], SatLas [83], and GFM [25]. This is particularly noteworthy as hierarchical transformers generally excel at generating multiscale features essential for object detection. Moreover, our SelectiveMAE with a ViT-B backbone outperforms SatMAE [30] and ScaleMAE [33], which use a larger ViT-L backbone, and matches the advanced RVSA [18, 34] that employs specialized window attention. When using the larger ViT-L backbone, SelectiveMAE shows significantly enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
@@ 6 Conclusion
In this paper, we introduce RS-4M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous datasets in the RS field, RS-4M offers a significantly larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on RS-4M highlights its advantages in these tasks. To address the computational burden of MIM training on large-scale RS datasets, we propose SelectiveMAE, an efficient MIM method that dynamically encodes and reconstructs patch tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that only 40% of RS image patches are needed to train a comparable MIM model. Experiments show that SelectiveMAE not only achieves over 2×2$\times$2 × speedup in pre-training but also enhances performance in various downstream tasks.
Limitations and Discussion. Although the RS-4M dataset surpasses existing RS datasets by more than four times, its size can be further increased by collecting and filtering additional RS data. Additionally, its diversity can be enhanced by incorporating data from various types of sensors.
    A: The dataset was created by:
    - Fengxiang Wang (National University of Defense Technology),
    - Hongzhen Wang (Tsinghua University),
    - Di Wang (Wuhan University),
    - Zonghao Guo (University of Chinese Academic of Sciences),
    - Zhenyu Zhong (Nankai University),
    - Long Lan (National University of Defense Technology),
    - Jing Zhang (The University of Sydney),
    - Zhiyuan Liu (Tsinghua University),
    - Maosong Sun (Tsinghua University).
  3. 3.
    “Who funded the creation of the dataset?”
    A: The dataset creation was funded by the affiliations of the authors involved in this work.
7.5.2 Composition
Most of the questions in this section are intended to provide dataset consumers with the information they need to make informed decisions about using the dataset for their chosen tasks. Some of the questions are designed to elicit information about compliance with the EU’s General Data Protection Regulation (GDPR) or comparable regulations in other jurisdictions. Questions that apply only to datasets that relate to people are grouped together at the end of the section. We recommend taking a broad interpretation of whether a dataset relates to people. For example, any dataset containing text that was written by people relates to people.
  1. 1.
    “What do the instances that comprise our datasets represent (e.g., documents, photos, people, countries)?”
    A: The dataset primarily comprises visible light remote sensing images captured by satellites. All datasets utilized in RS-4M are publicly accessible and nonprofit.
  2. 2.
    “How many instances are there in total (of each type, if appropriate)?”
    A: RS-4M contains 4 million remote sensing image instances captured by satellites.
  3. 3.
    “Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?”
    A: Yes, our dataset contains all possible instances that have been collected so far.
  4. 4.
    “Is there a label or target associated with each instance?”
    A: No, our dataset is intended for self-supervised learning. Therefore, each instance is an individual remote sensing image and does not contain annotations.
  5. 5.
    “Is any information missing from individual instances?”
    A: No.
  6. 6.
    “Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?”
    A: Yes, the relationship between individual instances is explicit.
  7. 7.
    “Are there recommended data splits (e.g., training, development/validation, testing)?”
    A: Yes, the entire dataset is intended for self-supervised methods, and we recommend using the whole dataset for self-supervised learning research.
  8. 8.
    “Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?”
    A: Yes, our dataset relies on many publicly available remote sensing datasets, which we have detailed in the main text.
  9. 9.
    “Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)?”
    A: No, all data are clearly licensed.
  10. 10.
    “Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?”
    A: No.
7.5.3 Collection Process
In addition to the goals outlined in the previous section, the questions in this section are designed to elicit information that may help researchers and practitioners create alternative datasets with similar characteristics. Again, questions that apply only to datasets that relate to people are grouped together at the end of the section.
  1. 1.
    “How was the data associated with each instance acquired?”
    A: Please refer to the details listed in the main text Sec. 3.
  2. 2.
    “What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?”
    A: Please refer to the details listed in the main text Sec. 3.
  3. 3.
    “If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?”
    A: Please refer to the details listed in the main text Sec. 3.
7.5.4 Preprocessing, Cleaning, and Labeling
The questions in this section are intended to provide dataset consumers with the information they need to determine whether the “raw” data has been processed in ways that are compatible with their chosen tasks. For example, text that has been converted into a “bag-of-words" is not suitable for tasks involving word order.
  1. 1.
    “Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?”
    A: Yes, we preprocessed and cleaned data in our dataset.
  2. 2.
    “Was the ‘raw’ data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?”
    A: Yes, raw data is accessible.
  3. 3.
    “Is the software that was used to preprocess/clean/label the data available?”
    A: Yes, the necessary software used to preprocess and clean the data is publicly available.
7.5.5 Uses
The questions in this section are intended to encourage dataset creators to reflect on tasks for which the dataset should and should not be used. By explicitly highlighting these tasks, dataset creators can help dataset consumers make informed decisions, thereby avoiding potential risks or harms.
  1. 1.
    “Has the dataset been used for any tasks already?”
    A: No.
  2. 2.
    “Is there a repository that links to any or all papers or systems that use the dataset?”
    A: Yes, we provide such links in our GitHub repository.
  3. 3.
    “What (other) tasks could the dataset be used for?”
    A: The dataset could be used for training the remote sensing foundation models with the self-supervised learning method.
  4. 4.
    “Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?”
    A: N/A.
  5. 5.
    “Are there tasks for which the dataset should not be used?”
    A: N/A.
7.5.6 Distribution
Dataset creators should provide answers to these questions prior to distributing the dataset either internally within the entity on behalf of which the dataset was created or externally to third parties.
  1. 1.
    “Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?”
    A: No.
  2. 2.
    “How will the dataset be distributed (e.g., tarball on website, API, GitHub)?”
    A: Very likely to be distributed by website, API, and GitHub repository.
  3. 3.
    “When will the dataset be distributed?”
    A: The datasets are publicly accessible.
  4. 4.
    “Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?”
    A: Yes, the dataset is under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License.
  5. 5.
    “Have any third parties imposed IP-based or other restrictions on the data associated with the instances?”
    A: No.
  6. 6.
    “Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?”
    A: No.
7.5.7 Maintenance
As with the questions in the previous section, dataset creators should provide answers to these questions prior to distributing the dataset. The questions in this section are intended to encourage dataset creators to plan for dataset maintenance and communicate this plan to dataset consumers.
  1. 1.
    “Who will be supporting/hosting/maintaining the dataset?”
    A: The authors of this work serve to support, host, and maintain the datasets.
  2. 2.
    “How can the owner/curator/manager of the dataset be contacted (e.g., email address)?”
    A: The curators can be contacted via the email addresses listed on our webpage [^4] https://github.com/Fengxiang23/SelectiveMAE..
  3. 3.
    “Is there an erratum?”
    A: There is no explicit erratum; updates and known errors will be specified in future versions.
  4. 4.
    “Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?”
    A: Yes, for the current version. Future updates (if any) will be posted on the dataset website.
  5. 5.
    “Will older versions of the dataset continue to be supported/hosted/maintained?”
    A: Yes. This is the first version of the release; future updates will be posted and older versions will be replaced.
  6. 6.
    “If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?”
    A: Yes, we provide detailed instructions for future extensions.
7.6 Limitation and Potential Societal Impact
In this section, we elaborate on the limitations and potential societal impact of this work.
7.6.1 Potential Limitations
While RS-4M provides a comprehensive benchmark for training the remote sensing foundation models with self-supervised learning methods, there are several limitations to consider:
  • Scope of Sensors: Although our benchmark includes 4 million visible light remote sensing images, it may not cover all possible real-world scenarios. There could be additional sensor data, like multispectral data that were not included in this study, potentially limiting the generalizability of our findings.
  • Model and Dataset Diversity: While our dataset is primarily focused on adapting to downstream tasks like detection and segmentation, there is undeniably less data available for these tasks compared to general scene classification. In the future, we should collect more data that is better suited for various downstream tasks.
  • Computation and Resource Requirements: Pre-training on the extensive RS-4M dataset, comprising 4 million images, demands substantial computational resources, despite our introduction of the efficient SelectiveMAE method. This may limit access to the benchmark for research groups lacking ample computational power.
7.6.2 Potential Negative Societal Impact
While the development of remote sensing foundation models with self-supervised learning methods has the potential to significantly advance remote sensing downstream tasks, there are potential negative societal impacts that must be considered:
  • Safety Risks: Our benchmark aims to improve remote sensing foundation models, but relying too heavily on these models may breed overconfidence in autonomous systems. It’s crucial to deploy these systems with adequate safety measures and human oversight to uphold public safety.
  • Environmental Impact: Training and evaluating remote sensing foundation models demand substantial computational resources, leading to a notable environmental footprint. Encouraging the adoption of energy-efficient algorithms (i.e., the proposed SelectiveMAE) and sustainable computing practices is crucial to reduce the environmental impact of this research.
  • Bias and Fairness: The performance of foundational remote sensing models can vary across different environments and conditions, potentially introducing biases in downstream tasks within the remote sensing domain. It is essential to train and evaluate these models on diverse datasets to ensure fairness and avoid discriminatory outcomes.

### v3

@@ 4.2.1 Fine-tuning Performance on Downstream Tasks
We further evaluated SelectiveMAE with the OpticalRS-4M dataset across three key downstream tasks: scene classification, object detection, and semantic segmentation. In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[54], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models. These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning-based approaches [56]. Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE). More details can be found in the Appendix.
Scene Classification. We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders. We leverage two scene classification datasets: AID [85] and RESISC-45 [86]. Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code. Evaluation is based on overall accuracy (OA). Table 6 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline. Besides, it outperforms RVSA and others when scaled to ViT-L. This demonstrates that, with our SelectiveMAE training method on the OpticalRS-4M dataset, the model can learn strong feature representations and scale efficiently.
@@
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [87] and its improved variant DIOR-R for oriented object detection [88], both comprising RGB images. Following the methodologies of previous works [18, 37], we maintained consistent experimental setups, employing Faster-RCNN [89] and Oriented-RCNN [90] as detectors for each dataset. Detailed experimental configurations are available in the accompanying code. The results are summarized in Table 6. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
@@ 5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. Built upon it, we first curated OpticalRS-4M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-4M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-4M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that only 40% RS image patches are needed to train a comparable MIM model. Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE achieves over 2×2$\times$2 × speedup, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.